Perform the conversions between energy units.
2185 kcal
7.29×106
1
Expert's answer
2019-10-17T07:30:05-0400

"1 \ Kcal=4184 \ joules."

"1 \ J=\frac{1}{4184}\ KCal"

"2185 \ KCal=2185 \times 4184\ J=9,142,040\ J\approx9.142\times 10^{6} \ J"

"7.29 \times 10^6J=\frac{7.29\times 10^6}{4184} \ KCal=1.742 \times10^3 KCal"